A coin box has $3.25 consisting of nickels and dimes. The number of dimes is 5 less than twice the number of nickels. How many of each kind are in the box?
Answer by mananth(16946)

nickels ---------x
dimes-----------(2x-5)
Value= 325

5x+20x -50=325
25x-50=325
25x=325+50
25x=375
/25
x=15 nickels
Dimes = 2x-5= 25